Find the value of k so that the line has slope m:
6x + ky = 10 and m = -2

Answer by nerdybill(7384)

Find the value of k so that the line has slope m:
6x + ky = 10 and m = -2
.
Start by rearranging:
6x + ky = 10
so that it is in "slope-intercept" form:
y = mx + b
.
6x + ky = 10
ky = -6x + 10
y = (-6/k)x + 10/k
Set the slope above to -2 and solve for k:
-6/k = -2
-6 = -2k
6 = 2k
3 = k